### v1

@@ 4 Fejér* and quasi-Fejér sequences
$\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\frac{1}{2}\left(\lim\left\|x^{j_{k}}\right\|^{2}-\lim\left\|x^{\ell_{k}}\right\|^{2}\right).$
@@
$M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\},$
@@
Item (iii): In view of item (ii), if $\operatorname{cl}\left({\operatorname{aff}(M)}\right)=\mathcal{H}$ then, for all $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, there exists an $\alpha\in\mathds{R}$ such that for all $y\in\mathcal{H}$ it holds that $\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha$. Consequently, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ reduces to a singleton. Due to Proposition 2(i), $\left(x^{k}\right)_{k\in\mathds{N}}$ lies in a weakly compact set and therefore converges weakly.
@@
$\left\|x^{k}-y\right\|^{2}=\left\|x^{k}-\bar{x}\right\|^{2}+2\left\langle x^{k}-\bar{x},\bar{x}-y\right\rangle+\|\bar{x}-y\|^{2}.$
Using the hypothesis and Proposition 2 on the right-hand side of the above equation, we conclude the convergence of $\left(\left\|x^{k}-y\right\|\right)_{k\in\mathds{N}}$.
Theorem 4.2
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then, $\left(x^{k}\right)_{k\in\mathds{N}}$ converges weakly to a point in $M$ if, and only if, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$.
Proof
It holds trivially that Fejér* monotonicity together with weakly convergence implies that $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$. On the other hand, suppose that $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$ and take $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$. Thus, $w^{\prime},w^{\prime\prime}\in M$. Proposition 3(ii) affirms that $\left\langle w^{\prime},w^{\prime}-w^{\prime\prime}\right\rangle=\left\langle w^{\prime\prime},w^{\prime}-w^{\prime\prime}\right\rangle$ and hence, $w^{\prime}=w^{\prime\prime}$. The boundedness of $\left(x^{k}\right)_{k\in\mathds{N}}$, guaranteed by Proposition 1, completes the proof. ∎
@@
Proposition 4
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then,
@@
    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y-\frac{w^{\prime}+w^{\prime\prime}}{2},w^{\prime}-w^{\prime\prime}\right\rangle=0\right\};$
@@
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}-\rho\left\|x^{k+1}-x^{k}\right\|+\varepsilon_{k}.$
Proof
It follows from Theorem 4.1 together with Proposition 3.7 in Combettes:2001a . ∎
Theorem 4.3
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then, the following statements are equivalent:
@@
Proof
This is a direct result of combining Theorem 4.1 and Theorem 3.11 in Combettes:2001a . ∎
We emphasize that new characterizations of Fejér* monotonicity could be revealed with further analysis, suggesting a direction for future research.
@@ 5 Concluding remarks
In this work we further expand the understanding of Fejér* monotonicity. We provide a characterization of Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull of the underlying set. We provide an insightful example that shows that Fejér* monotonicity not always can be extended to the closure of the convex hull of the underlying set. It was also established connections between Fejér* monotonicity and quasi-Fejér monotonicity. The results presented in this paper could shed a light on the analysis of the (weak) convergence of algorithms that comply with Fejér* monotonicity.
Acknowledgements.
RB was partially supported by Brazilian agency Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro (Grant E-26/201.345/2021); YBC was partially supported by the USA agency National Science Foundation (Grant DMS-2307328); AAR was partially supported by Brazilian agency Conselho Nacional de Desenvolvimento Científico e Tecnológico – CNPq (Grant 307987/2023-0); LRS was also partially funded by CNPq (Grant 310571/2023-5).

### v2

@@ 4 Fejér* and quasi-Fejér sequences
$\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\frac{1}{2}\left(\lim_{k\to\infty}\left\|x^{j_{k}}\right\|^{2}-\lim_{k\to\infty}\left\|x^{\ell_{k}}\right\|^{2}\right).$ (39)
@@
$M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\},$ (40)
@@
Item (iii): In view of item (ii), if $\operatorname{cl}\left({\operatorname{aff}(M)}\right)=\mathcal{H}$ then, for all $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, there exists an $\alpha\in\mathds{R}$ such that for all $y\in\mathcal{H}$ it holds that $\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha$. Consequently, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ reduces to a singleton. Due to Proposition˜3.8(i), $\left(x^{k}\right)_{k\in\mathds{N}}$ lies in a weakly compact set and therefore converges weakly.
@@
$\left\|x^{k}-y\right\|^{2}=\left\|x^{k}-\bar{x}\right\|^{2}+2\left\langle x^{k}-\bar{x},\bar{x}-y\right\rangle+\|\bar{x}-y\|^{2}.$ (41)
Using the hypothesis and Proposition˜3.8 on the right-hand side of the above equation, we conclude the convergence of $\left(\left\|x^{k}-y\right\|\right)_{k\in\mathds{N}}$.
Theorem 4.5.
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a sequence in $\mathcal{H}$ and let $M$ be a nonempty subset of $\mathcal{H}$. Suppose that $\left(x^{k}\right)_{k\in\mathds{N}}$ is Fejer* monotone with respect to $M$. Then, $\left(x^{k}\right)_{k\in\mathds{N}}$ converges weakly to a point in $M$ if, and only if, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$.
Proof.
The result follows directly from Theorem˜4.2 and [Combettes:2001a, Thm. 3.8]. ∎
@@
Proposition 4.6.
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a sequence in $\mathcal{H}$ and let $M$ be a nonempty subset of $\mathcal{H}$. Suppose that $\left(x^{k}\right)_{k\in\mathds{N}}$ is Fejer* monotone with respect to $M$. Then,
@@
    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y-\frac{w^{\prime}+w^{\prime\prime}}{2},w^{\prime}-w^{\prime\prime}\right\rangle=0\right\};$ (42)
@@
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}-\rho\left\|x^{k+1}-x^{k}\right\|+\varepsilon_{k}.$ (43)
Proof.
It follows from Theorem˜4.2 together with Proposition 3.9 in [Combettes:2001a]. ∎
Theorem 4.7.
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a sequence in $\mathcal{H}$ and let with $M$ be a nonempty subset of $\mathcal{H}$. Suppose that $\left(x^{k}\right)_{k\in\mathds{N}}$ is Fejer* monotone with respect to $M$. Then, the following statements are equivalent:
@@
Proof.
This is a direct result of combining Theorem˜4.2 and Theorem 3.11 in [Combettes:2001a].
@@ 5 Concluding remarks
In this work, we further expanded the understanding of Fejér* monotonicity. We provided a characterization of Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull of the underlying set. The insightful Example˜3.3 illustrates that Fejér* monotonicity cannot, in general, be extended to the closure of the convex hull of the underlying set. We also established connections between Fejér* monotonicity and quasi-Fejér monotonicity, highlighting that Fejér* monotonicity lies outside the classical Fejér and quasi-Fejér of Type II frameworks. Fejér* monotonicity notion has already been used to establish convergence (usually finite) of algorithms in [Iusem:1986a, fukushimaFinitelyConvergentAlgorithm1982, iusemFinitelyConvergentIterative1987, DePierro:1988a, Kolobov:2022] and, more recently, was formally defined in [Behling:2024c]. The results presented in this paper could be used to shed light on the analysis of the (weak) convergence of algorithms that comply with Fejér* monotonicity, mainly to aid to prove finite convergence results.
[…]
Acknowledgements The authors would like to thank the anonymous reviewers for their careful reading of our manuscript and their insightful comments and suggestions that helped us to improve the quality of this paper. In particular, the discussions presentend in Examples˜3.5, 3.6, and 3.7 were raised by one of the reviewers. RB was partially supported by Brazilian agency Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro (Grant E-26/201.345/2021); YBC was partially supported by the USA agency National Science Foundation (Grant DMS-2307328); AAR was partially supported by Brazilian agency Conselho Nacional de Desenvolvimento Científico e Tecnológico – CNPq (Grant 307987/2023-0); LRS was also partially funded by Brazilian agencies CNPq (Grants 310571/2023-5 and 407147/2023-3) and Fundação de Amparo ao Estado de Santa Catatrina – FAPESC (Edital 21/2024, Grant 2024TR002238).
[…]
AI use declaration During the preparation of this work, the authors used Google Gemini (Large Language Model) for language editing, generating code, and refining the clarity of the text. The authors reviewed and verified all AI-generated content and take full responsibility for the accuracy and integrity of the publication.